klmjnmkj:

Determine the direction in which the graph of the following parabola opens. f(x)=(x−5)^2−1

klmjnmkj:

1 attachment
klmjnmkj:

is it B

klmjnmkj:

@Florisalreadytaken

Florisalreadytaken:

firstly, why did you chose down?

klmjnmkj:

necause it looks like it goes don

Florisalreadytaken:

how?

klmjnmkj:

1 attachment
Florisalreadytaken:

ah you did it through graphing -- does it look like it opens downwards tho?

Florisalreadytaken:

we can actually see that the parabola opens UPWARDS, so thatd be your answer

klmjnmkj:

Determine the axis of symmetry of the graph of the following parabola. @Florisalreadytaken

klmjnmkj:

(5, -1)?

klmjnmkj:

1 attachment
klmjnmkj:

@snowflake0531

Florisalreadytaken:

is it for the same parabola? or a different one?

klmjnmkj:

its the same parabola

Florisalreadytaken:

ah ok -- so the axis of symetry would be \[ x= -\frac{b}{2a} \]

klmjnmkj:

how to get b and a

Florisalreadytaken:

we teached you that earlier lol -- imma give you some time to remember (since you got time)

Florisalreadytaken:

snowflake do not tell him

Florisalreadytaken:

i want him to solve it

snowflake0531:

We first have to factor this out, we have ot get this to standard form \[(x-5)^2 -1\] what does this equal, remember that \[(a-b)^2 =a^2 -2ab+b^2\]

klmjnmkj:

4 and 6?

snowflake0531:

yes

klmjnmkj:

x^2-10x+24

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @klmjnmkj x^2-10x+24 \(\color{#0cbb34}{\text{End of Quote}}\) Yes so from this we get (x-6)(x-4) which gets us x=6 and x=4, as you've already got above

klmjnmkj:

alright

klmjnmkj:

whats next

klmjnmkj:

umm now

snowflake0531:

OH i didnt' even read your question, i thought youw ere just solving for the zeroes

snowflake0531:

So you had that x^2 -10x +24 a=1 b=-10 put this into -b/2a

klmjnmkj:

-10/2 1

klmjnmkj:

i divid?

snowflake0531:

-10/2 simplifies to what

klmjnmkj:

-5

Florisalreadytaken:

do not forget we had the - at the front

Florisalreadytaken:

so - times - = +

Florisalreadytaken:

thus x=5

klmjnmkj:

alright

klmjnmkj:

the ansers 5?

Florisalreadytaken:

answer* yes